1. The pdf of Y is f(y) = k(y^3) I[2,3](y) where k is a constant, real number.

(a) (5 points) Calculate the value of k so that f(y) is a valid pdf.

(b) (5 points) Find E(Y ). Hint: write k as a fraction so you don’t have to round until the end.

(c) (5 points) Find the 50th percentile of Y .

(d) (2 points) Briefly (1-2 sentences) explain why your answers to parts (b) and (c) make sense.
Answer #1

पूर to delys = ky² Itazy de july) = {hy elke as ļ ky dy = 1 kg ů 3-2]=1 3) 658 = 13) 6:4 b) EXT = (39-2°) 3 meses tall = 84